what is the slope in 2y+4 = x+ 3
Answer by stanbon(75887)

what is the slope in 2y+4 = x+ 3
----------
2y = x - 1
---
y = (1/2)x - (1/2)
-----------------------
slope = 1/2
==========================
Cheers,
Stan H.